# Dynamic phase transition into a mixed-CDW state in 1T-TaS2 via a thermal quench

### v2

@@
A. de la Torre a.delatorreduran@northeastern.edu Quantum Materials and Sensing Institute, Northeastern University, Burlington, Massachusetts, 01803, USA Department of Physics, Northeastern University, Boston, Massachusetts, 02115, USA Department of Physics, Brown University, Providence, Rhode Island 02912, United States    Q. Wang Department of Physics, Brown University, Providence, Rhode Island 02912, United States    B. Campbell Department of Physics and Astronomy, University of New Hampshire, Durham, NH 03824.    J. V. Riffle Department of Physics and Astronomy, University of New Hampshire, Durham, NH 03824.    D. Balasundaram Department of Physics and Astronomy, George Mason University, Fairfax, VA, 22030 USA Quantum Science and Engineering Center, George Mason University, Fairfax, VA, 22030 USA    P.M.Vora Department of Physics and Astronomy, George Mason University, Fairfax, VA, 22030 USA Quantum Science and Engineering Center, George Mason University, Fairfax, VA, 22030 USA    J.P.C. Ruff Cornell High Energy Synchrotron Source, Cornell University, Ithaca, NY, 14853, USA    S. M. Hollen Department of Physics and Astronomy, University of New Hampshire, Durham, NH 03824.    K. W. Plumb kemp˙plumb@brown.edu Department of Physics, Brown University, Providence, Rhode Island 02912, United States
(July 12, 2024)
@@
Ultrafast light-matter interaction has emerged as a new mechanism to exert control over the macroscopic properties of quantum materials toward novel functionality. To date, technological applications of these non-thermal phases are limited by their ultrashort lifetimes and low-ordering temperatures. Among the most studied photoinduced metastable phases for their technological promise is the hidden metallic charge density wave (H-CDW) in the model correlated CDW compound $1T$-TaS2. Despite active study and engineering, the nature of the photoinduced H-CDW remains the subject of debate and is only accessible at cryogenic temperatures. Here, we stabilize the H-CDW phase at thermal equilibrium up to near-room temperature by accessing an intermediate mixed CDW order regime via thermal quenching. Using x-ray high dynamic range reciprocal space mapping (HDRM) and scanning tunneling spectroscopy (STS), we reveal the coexistence of commensurate (C) CDW and H-CDW domains below 180 K during cooling and below 210 K during warming. Our findings show that each order parameter breaks basal plane mirror symmetry with different chiral orientations and induces out-of-plane unit cell tripling in the H-CDW phase. Despite metallic domain walls and a finite density of states at zero bias observed via STS, bulk resistance remains insulating due to CDW stacking disorder. This study establishes the H-CDW as a thermally stable phase and introduces a new mechanism for switchable metallic behavior in thin flakes of $1T$-TaS2 and similar materials with competing order phases.
@@ I Introduction
The correlated van der Waal material $1T$-TaS2 displays a series of first-order phase transitions associated with the onset of CDW orders with different degrees of commensurability Thompson et al. (1971); Wilson et al. (1975); Fazekas and Tosatti (1979). First, at $T=550$ K, a mirror-symmetry preserving incommensurate (I) CDW onsets from a metallic high-temperature state Sipos et al. (2008). At $T_{\mathrm{I-NC}}$$\approx 350$ K, $1T$-TaS2 enters into a nearly commensurate (NC) CDW phase. The atomic rearrangement associated with the incommensurate NC-CDW (in-plane ordering wavevector $q_{||}\!=\!0.284$ r.l.u.) leads to disconnected regions displaying a $\sqrt{13}\times\sqrt{13}$ structural reconstruction in which twelve Ta atoms move towards a central one forming a star-of-David (SOD) pattern Spijkerman et al. (1997). The NC-CDW breaks the mirror plane symmetries of the lattice basal plane and spontaneously selects one of two chiral domains rotated from the $a$-axis by $\phi\approx 12^{\circ}$. Then, at $T_{\mathrm{NC-C}}\approx 175$ K in cooling and $T_{\mathrm{NC-C}}\approx 225$ K in warming, a 3D CDW order emerges from the NC-CDW phase. This low-temperature phase is commensurate (C) with the atomic lattice in-plane ($\phi\!=\!13.9^{\circ}$ and $q_{||}\!=\!0.277$ r.l.u.) with the in-plane C-CDW atomic reconstruction inheriting the chiral structure from the NC-CDW. On the other hand, in the out-of-plane direction, the C-CDW displays a large stacking disorder between dimerized CDW planes Thompson et al. (1971); Wilson et al. (1975).
@@
In general, experimental probes spanning a large range of lengthscales find that $1T$-TaS2 samples lack domain twinning that is expected across a first-order phase transition Zong et al. (2018); Fichera et al. (2020); Luo et al. (2021); Yang et al. (2022); Campbell et al. (2024). This reflects the kinetics of the first cooling across the only first-order phase transition in $1T$-TaS2 that breaks in-plane mirror symmetries, $T_{\mathrm{I-NC}}$ Sept and Tuszyński (1994); Binder (1987). During synthesis, the $1T$ phase is stabilized by quenching to $T\approx 273$ K across the polytype transition, $T_{\rm{1T-2H}}\!=\!600$ K Thompson et al. (1971). During this initial quench, the cooling rate across $T_{\mathrm{I-NC}}$ is effectively slow when compared to the intrinsic timescales of the electronic ($\approx$ fs) and atomic ($\approx$ ps) degrees of freedom and enables the nucleation of large chiral CDW domains with lateral sizes comparable to the correlation length of the order parameter parameter Oxtoby (1998). The chiral domain structure of the C-CDW in $1T$-TaS2 is stable to cooling through $T_{\mathrm{NC-C}}$ Zong et al. (2018); Fichera et al. (2020); Luo et al. (2021); Yang et al. (2022); Campbell et al. (2024) with suggestions that strain fields and defects can pin a single chirality Zong et al. (2018); Salzmann et al. (2023a). Still, chirality reversal of single domain samples has been demonstrated by repeated cleaving of a bulk sample Fichera et al. (2020) or by cycling through $T_{\mathrm{I-NC}}$ Zhao et al. (2023). On the other hand, multi-chiral domain samples can be engineered in thin exfoliated flakes by heating and quenching above $T_{\rm{1T-2H}}$ Sung et al. (2022); Husremović et al. (2023) or upon irradiation with 800 nm laser pulses with fluence $F>7$ mJ/cm2 Zong et al. (2018). Once created, chiral domain walls are long-lived at low temperatures, but a dominant chiral domain of dimensions comparable to the sample size can be restored by heating above $T_{\mathrm{I-NC}}$ and adiabatically cooling through the mirror symmetry-breaking IC-NC transition Sung et al. (2022); Husremović et al. (2023); Zong et al. (2018).Thus, $T_{\mathrm{I-NC}}$ emerges as the relevant energy scale for chiral CDW domain nucleation in $1T$-TaS2 as the initial mirror-symmetry breaking transition in the CDW phase diagram Thompson et al. (1971). Moreover, x-ray pair distribution function has revealed that the in-plane Ta average displacement from the mean crystallographic position, a measure of the CDW formation, remains constant across the NC-C CDW transition. At the same time, it dramatically changes across $T_{\mathrm{I-NC}}$ Bozin et al. (2023).
@@
An additional low-temperature metastable metallic CDW phase, the so-called hidden H-CDW phase, has been accessed in $1T$-TaS2 from the insulating C-CDW state via a single femtosecond laser pulse of fluence larger than a critical value ($F_{c}\!>\!1$ mJ/cm2) Stojchevska et al. (2014); Stahl et al. (2020); Vaskivskyi et al. (2015) or by exciting $1T$-TaS2 with a ns voltage pulse Hollander et al. (2015); Ma et al. (2016); Cho et al. (2016); Venturini et al. (2022); Tsen et al. (2015). The H-CDW phase locally resembles the SOD arrangement but breaks the large C-CDW domains into many nanoscale domains and has a metallic density of states Gerasimenko et al. (2019); Ma et al. (2016); Maklar et al. (2023) These domains maintain the chirality of the C-CDW Stahl et al. (2020) and are related by a relative translation of the SOD pattern by a finite number of atomic lattice vectors Ma et al. (2016); Gerasimenko et al. (2019); Ravnik et al. (2021). Moreover, slight differences between the displacement of the Ta atoms between the H-CDW and the C-CDW distinguish the two. The H-CDW phases result in CDW Bragg peaks rotated by $\phi\!=\!11.9^{\circ}$ from the $a^{*}$-axis with a larger basal plane ordering wavevector, $q_{||}\!=\!0.284$ r.l.u.. The CDW plane stacking also rearranges towards a $l\!=\!1/3$ periodicity along the c-axis Stahl et al. (2020). We note that a similar metallic mosaic state can also be induced by cooling thin exfoliated samples ($d<24$ nm) across $T_{\mathrm{I-NC}}$ with a cooling rate faster than $0.2$ K/s Yoshida et al. (2015) or by intrinsic strain and defect effects Mutka et al. (1981); Zhang et al. (2022); Salzmann et al. (2023b).
@@
Despite much study, the origin of the H-CDW state remains controversial, with photodoping Sun et al. (2018); Gerasimenko et al. (2019), incoherent thermal effects Ma et al. (2016); Gao et al. (2022) or electron-phonon relaxation bottlenecks Maklar et al. (2023); Mante et al. (2022) all possible explanations. Nonetheless, the H-CDW has attracted much interest, not only as an example of the non-thermal control of quantum materials de la Torre et al. (2021) but also as a platform for new functional devices Mihailovic et al. (2021); Mraz et al. (2022); Devidas et al. (2024). However, the H-CDW has only been stabilized for ultrashort periods (${}100\mu$s) at temperatures above $T\!=\!50$ K Stojchevska et al. (2014); Ravnik et al. (2018); Maklar et al. (2023) preventing its deployment in efficient and fast memory devices towards novel technological applications Basov et al. (2017).
@@
Here, we demonstrate that the H-CDW phase exists at thermal equilibrium in $1T$-TaS2 and extend its thermal stability to near-room temperature by accessing a new long-lived mixed-CDW state via a reversible thermal quench across $T_{\mathrm{I-NC}}$. We deploy x-ray high dynamic range reciprocal space mapping (HDRM) to unequivocally demonstrate the co-existence of the C-CDW and H-CDW states in thermally quenched samples. Each CDW order parameter is linked to a mirror-symmetry breaking chiral domain. The mixed-CDW state emerges from a multi-domain NC-CDW state below $T^{CD}=180$ K in cooling and $T^{WU}=210$ K in warming. The mixed-CDW remains stable up to $T_{\mathrm{I-NC}}$. Scanning tunneling spectroscopy (STS) shows that the mixed-CDW is characterized by a semi-metallic in-plane density of states. At the same time, bulk resistivity remains insulating due to out-of-plane CDW stacking disorder in mixed CDW bulk samples. Our results demonstrate a new state in the phase diagram of $1T$-TaS2, enabling a new mechanism to engineer a semimetallic density of states at near room temperature.
@@ II Results
Accessing a novel mixed-CDW state by quenching through the mirror symmetry breaking transition. In Figure 1 a, we show a characteristic [HK0] plane reciprocal space map for the C-CDW phase of $1T$-TaS2, collected after a standard cooling cycle. Red circular markers highlight the CDW reflections at $q^{C}_{1}\!=\!(\sigma^{C}_{1},\sigma^{C}_{2},l)$ and $q^{TH}_{2}\!=\!(-\sigma^{C}_{2},\sigma^{C}_{1}+\sigma^{C}_{2}+,l)$ with $\sigma^{C}_{1}\!=\!3/13$ r.l.u and $\sigma^{C}_{2}\!=\!1/13$ r.l.u Thompson et al. (1971); Wilson et al. (1975), corresponding to a single C-CDW chiral domain. Additional non-indexable reflections in the reciprocal space map are attributed to a second small crystallite in our sample. Identical reciprocal space maps were obtained across many cooling cycles and samples. However, we found that the C-CDW domain structure depends on the cooling rate, and a fast thermal quench across $T_{\mathrm{I-NC}}$ can consistently produce a multi-domain CDW. Fig. 1 b shows the [HK0] reciprocal space map after an in-situ quench in our bulk sample across $T_{\mathrm{I-NC}}$ down to $T\!=\!140$ K at a rate of $0.2$ K/ms near $T_{\mathrm{I-NC}}$ (See Methods for a description of the in-situ quench method and calculation of quenching rates using heat diffusion calculations). We found that thermal quenching consistently resulted in a new set of superlattice reflections of opposite chirality to the original set of satellite peaks ($\phi\!=\!-12.3^{\circ}$) [Fig. 1 b]. The superior reciprocal space resolution of synchrotron single crystal HDRM enables us to resolve differences in the wavevector of the two sets of CDW Bragg peaks with different chirality. The in-plane wavevector magnitude $|q_{||}|\!=\!0.284(3)~{}\AA^{-1}$ of the quenched-induced satellite peaks is larger than that of the C-CDW ($|q_{||}|\!=\!0.277(3)~{}\AA^{-1}$) and agrees with that of the light-stabilized H-CDW state Stahl et al. (2020). This is shown in Fig. 1 c, where we display the integrated intensity across the C-CDW and H-CDW Bragg peaks (dashed lines in Fig. 1 a and b) as a function of $|q_{||}|$. Strikingly, each order parameter preferentially selects a chirality orientation, with the C-CDW domains of the mixed-CDW phase maintaining that of the as-grown samples.
@@
We also find that the c-axis periodicity differs between the C-CDW and H-CDW-like domains. Figure 1 d and e show [H0L] reciprocal space maps centered at H = 1 r.l.u and L = 0 r.l.u before and after the thermal quench, respectively. The as-grown sample displays a broad two-peak structure, with the center of mass at integer values of L. These short-range out-of-plane correlations are characteristic of the dimerization of CDW layers in the low-temperature C-CDW phase (Fig. 1 g) Tanda et al. (1984). On the other hand, the H-CDW-like satellite peaks shift towards $l\!=\!\pm 1/3$ indicating a collapse of the interlayer dimerization towards a triple layer CDW stacking (Fig. 1 h) as was observed for the light-induced H-CDW Stahl et al. (2020). The H-CDW-like satellite peaks are sharper than the C-CDW ones, reflecting a longer correlation length for the quench-induced domains $\xi\!=\!6(6)$ nm (Fig. 1 f). Our HDRM data in quenched samples demonstrates the coexistence of the C-CDW with a second CDW phase with the same order parameter as the H-CDW phase at much higher temperatures ($T=140$ K).
@@
The stability of the H-CDW phase at temperatures below $T_{\mathrm{I-NC}}$ in the mixed CDW phase enables us to ex-situ engineer the CDW mixed state at high temperatures (see Methods). We note that the mixed CDW state is stable over months, independent of the number of cleaves, and has been reproduced in quenched samples over multiple growth batches. Detailed temperature-dependent XRD measurements (See Methods) are shown in Fig. 2 f. All ex-situ quenched samples display twelve satellite reflections characteristic of the mixed-CDW state at low temperatures. Upon warming above $T^{WU}\!=\!210$ K, or cooling below $T^{CD}\!=\!195$ K, we observe a transition into an NC-CDW state with CDW Bragg peaks of two opposite chiralities. The existence of two chiral states in the NC-CDW is consistent with previous ultrafast electron diffraction experiments in $1T$-TaS2 driven with a low fluence optical pump reporting that the chirality of the low-temperature C-CDW domains is inherited from the symmetry-breaking IC-NC transition Haupt et al. (2016). We remark that this is distinct from the metastable H-CDW phase, which directly emerges from the C-CDW while maintaining the same chirality Stahl et al. (2020); Stojchevska et al. (2014); Gerasimenko et al. (2019). Detailed temperature dependence of the H-CDW-like peak and its NC-CDW precursor are shown in Fig. 2 h and i. C-CDW and H-CDW-like Bragg peaks show the same temperature dependence (see Extended data) suggesting the coupled nature of the two order parameters in the mixed-CDW phase. The asymmetrically narrower hysteretic loop of the first-order transition, when compared to as-grown samples, $T^{CD}\!=\!180$ K and $T^{WU}\!=\!210$ K, in conjunction with the narrower CDW Bragg peaks along L and the temperature dependence of the [0010] structural Brag peak shown in Fig. 2 c, suggests longer range interlayer ordering and a more 3D nature of the H-CDW phase.

### v3

@@
Alberto de la Torre a.delatorreduran@northeastern.edu Department of Physics, Northeastern University, Boston, Massachusetts, 02115, USA Quantum Materials and Sensing Institute, Northeastern University, Burlington, Massachusetts, 01803, USA Department of Physics, Brown University, Providence, Rhode Island 02912, United States    Qiaochu Wang Department of Physics, Brown University, Providence, Rhode Island 02912, United States    Yasamin Masoumi Department of Physics, Northeastern University, Boston, Massachusetts, 02115, USA    Benjamin Campbell Department of Physics and Astronomy, University of New Hampshire, Durham, NH 03824.    Jake V. Riffle Department of Physics and Astronomy, University of New Hampshire, Durham, NH 03824.    Dushyanthini Balasundaram Department of Physics and Astronomy, George Mason University, Fairfax, VA, 22030 USA Quantum Science and Engineering Center, George Mason University, Fairfax, VA, 22030 USA    Pattrick M.Vora Department of Physics and Astronomy, George Mason University, Fairfax, VA, 22030 USA Quantum Science and Engineering Center, George Mason University, Fairfax, VA, 22030 USA    Jacob P.C. Ruff Cornell High Energy Synchrotron Source, Cornell University, Ithaca, NY, 14853, USA    Gregory A. Fiete Department of Physics, Northeastern University, Boston, Massachusetts, 02115, USA Quantum Materials and Sensing Institute, Northeastern University, Burlington, Massachusetts, 01803, USA Department of Physics, Massachusetts Institute of Technology, Cambridge, MA 02139, USA    Shawna M. Hollen Department of Physics and Astronomy, University of New Hampshire, Durham, NH 03824.    Kemp W. Plumb kemp˙plumb@brown.edu Department of Physics, Brown University, Providence, Rhode Island 02912, United States
(January 8, 2025)
@@
Ultrafast light-matter interaction has emerged as a new mechanism to exert control over the macroscopic properties of quantum materials toward novel functionality. To date, technological applications of non-thermal phases are limited by their ultrashort lifetimes and low-ordering temperatures. The hidden metallic charge density wave (H-CDW) in $1T$-TaS2 is among the most studied photoinduced metastable phases because of its technological promise. However, despite active study, the nature of the photoinduced H-CDW remains the subject of debate and potential applications have been limited because it has so far only been stabilized at cryogenic temperatures. Here, we stabilize the H-CDW phase at thermal equilibrium by accessing a mixed CDW order regime via thermal quenching. Using x-ray high dynamic range reciprocal space mapping (HDRM) and scanning tunneling spectroscopy (STS), we reveal the coexistence of commensurate (C) CDW and H-CDW domains up to 210 K. Our findings show that each order parameter breaks basal plane mirror symmetry with different chiral orientations and induces out-of-plane unit cell tripling in the H-CDW phase. Despite metallic domain walls and a finite density of states at zero bias observed via STS, bulk resistance remains insulating due to CDW stacking disorder. By comparing our data to Landau-Ginzburg theory calculations of the free energy functional, our study establishes the H-CDW as a thermally stable phase and introduces a new mechanism for switchable metallic behavior in thin flakes of $1T$-TaS2 and similar materials with competing phases.
@@ I Introduction
The correlated van der Waals material $1T$-TaS2 displays a series of first-order phase transitions associated with the onset of CDW orders with different degrees of commensurability Thompson et al. (1971); Wilson et al. (1975); Fazekas and Tosatti (1979). Lowering from high temperatures, at $T=550$ K a mirror-symmetry preserving incommensurate (I) CDW onsets from a metallic state Sipos et al. (2008). At $T_{\mathrm{I-NC}}$$\approx 350$ K, $1T$-TaS2 enters into a nearly commensurate (NC) CDW phase. The atomic rearrangement associated with the incommensurate NC-CDW (in-plane ordering wavevector $q_{||}\!=\!0.284$ r.l.u.) leads to disconnected regions displaying a $\sqrt{13}\times\sqrt{13}$ spatial reconstruction in which twelve Ta atoms move towards a central one forming a star-of-David (SOD) pattern Spijkerman et al. (1997). The NC-CDW breaks the mirror symmetries of the lattice basal plane and spontaneously selects one of two chiral domains rotated from the $a$-axis by $\phi\approx 12^{\circ}$. Then, at $T_{\mathrm{NC-C}}\approx 175$ K in cooling and $T_{\mathrm{NC-C}}\approx 225$ K in warming, a 3D CDW order emerges from the NC-CDW phase. This low-temperature phase is commensurate (C) with the atomic lattice in-plane ($\phi\!=\!13.9^{\circ}$ and $q_{||}\!=\!0.277$ r.l.u.) with the in-plane C-CDW atomic reconstruction inheriting the chiral structure from the NC-CDW. On the other hand, in the out-of-plane direction, the C-CDW displays a large stacking disorder between dimerized CDW planes Thompson et al. (1971); Wilson et al. (1975).
@@
Experimental probes spanning a large range of length scales find that in the C-CDW phase of $1T$-TaS2 samples lack the chiral twin domains expected across the $T_{\mathrm{NC-C}}$ first-order phase transition Zong et al. (2018); Fichera et al. (2020); Luo et al. (2021); Yang et al. (2022); Campbell et al. (2024). The stabilization of a single chiral domain reflects the kinetics of the cooling across the only first-order phase transition in $1T$-TaS2 that breaks in-plane mirror symmetries, $T_{\mathrm{I-NC}}$ Sept and Tuszyński (1994); Binder (1987). During synthesis, the metastable $1T$ phase is stabilized by quenching from $T\approx 1000$ K to $T\approx 273$ K across the polytype transition, $T_{\rm{1T-2H}}\!=\!600$ K Thompson et al. (1971). During this initial quench, the cooling rate across $T_{\mathrm{I-NC}}$ is effectively slow when compared to the intrinsic timescales of the electronic ($\approx$ fs) and lattice ($\approx$ ps) degrees of freedom and enables the nucleation of large chiral CDW domains with lateral sizes comparable to the correlation length of the order parameter parameter Oxtoby (1998). The chiral mono-domain structure of the C-CDW in $1T$-TaS2 is stable to cooling through $T_{\mathrm{NC-C}}$ Zong et al. (2018); Fichera et al. (2020); Luo et al. (2021); Yang et al. (2022); Campbell et al. (2024) with suggestions that strain fields and defects can pin a single chirality Zong et al. (2018); Salzmann et al. (2023a). Still, chirality reversal of single domain samples has been demonstrated by repeated cleaving of a bulk sample Fichera et al. (2020) or by cycling through $T_{\mathrm{I-NC}}$ Zhao et al. (2023).
On the other hand, multi-chiral domain samples can be engineered in thin exfoliated flakes by heating and quenching above $T_{\rm{1T-2H}}$ Sung et al. (2022); Husremović et al. (2023) or upon irradiation with 800 nm laser pulses with fluence $F>7$ mJ/cm2 Zong et al. (2018). Once created, chiral domain walls are long-lived at low temperatures, but a dominant chiral domain of dimensions comparable to the sample size can be restored by heating above $T_{\mathrm{I-NC}}$ and adiabatically cooling through the mirror symmetry-breaking IC-NC transition Sung et al. (2022); Husremović et al. (2023); Zong et al. (2018). Thus, $T_{\mathrm{I-NC}}$ emerges as the relevant energy scale for chiral CDW domain nucleation in $1T$-TaS2 as the temperature at which the initial mirror-symmetry breaking transition in the CDW phase diagram Thompson et al. (1971) occurs and the largest specific heat divergence is observed Mañas-Valero et al. (2021). Moreover, x-ray pair distribution function measurements have revealed that the anisotropic Ta Debye-Waller factor, associated with CDW formations, remains constants across the NC-C CDW transition but dramatically changes across $T_{\mathrm{I-NC}}$ Bozin et al. (2023).
@@
An additional low-temperature metastable metallic CDW phase, the so-called hidden H-CDW phase, has been accessed in $1T$-TaS2 from the insulating C-CDW state via a single femtosecond laser pulse of fluence larger than a critical value ($F_{c}\!>\!1$ mJ/cm2) Stojchevska et al. (2014); Stahl et al. (2020); Vaskivskyi et al. (2015) or by exciting $1T$-TaS2 with a ns voltage pulse Hollander et al. (2015); Ma et al. (2016); Cho et al. (2016); Venturini et al. (2022); Tsen et al. (2015). The H-CDW phase locally resembles the SOD arrangement but breaks the large C-CDW domains into many nanoscale domains and has a metallic density of states Gerasimenko et al. (2019); Ma et al. (2016); Maklar et al. (2023). These domains maintain the chirality of the C-CDW Stahl et al. (2020) and are related by a relative translation of the SOD pattern by a finite number of atomic lattice vectors Ma et al. (2016); Gerasimenko et al. (2019); Ravnik et al. (2021). Slight differences between the displacement of the Ta atoms between the H-CDW and the C-CDW distinguish the two. The H-CDW phases result in CDW Bragg peaks rotated by $\phi\!=\!11.9^{\circ}$ from the $a^{*}$-axis with a larger basal plane ordering wavevector, $q_{||}\!=\!0.284$ r.l.u.. The CDW plane stacking also rearranges towards a $l\!=\!1/3$ periodicity along the c-axis Stahl et al. (2020). We note that a similar metallic mosaic state can also be induced by cooling thin exfoliated samples ($d<24$ nm) across $T_{\mathrm{I-NC}}$ with a cooling rate faster than $0.2$ K/s Yoshida et al. (2015) or by intrinsic strain and defect effects Mutka et al. (1981); Zhang et al. (2022); Salzmann et al. (2023b).
@@
Despite much study, the origin of the H-CDW state remains controversial, with photodoping Sun et al. (2018); Gerasimenko et al. (2019), incoherent thermal effects Ma et al. (2016); Gao et al. (2022) or electron-phonon relaxation bottlenecks Maklar et al. (2023); Mante et al. (2022) all possible explanations. Nonetheless, the H-CDW has attracted much interest, not only as an example of the non-thermal control of quantum materials de la Torre et al. (2021) but also as a platform for new functional devices Mihailovic et al. (2021); Mraz et al. (2022); Devidas et al. (2024). However, the H-CDW has only been stabilized for ultrashort periods ($\sim 100\,\,\mu$s) at temperatures above $T\!=\!50$ K Stojchevska et al. (2014); Ravnik et al. (2018); Maklar et al. (2023) preventing its deployment in efficient and fast memory devices towards novel technological applications Basov et al. (2017).
@@
Here, we demonstrate that the H-CDW phase exists in thermal equilibrium in $1T$-TaS2 and extend its thermal stability by a factor of three by accessing a new long-lived mixed-CDW state via a reversible thermal quench across $T_{\mathrm{I-NC}}$. We use x-ray high dynamic range reciprocal space mapping (HDRM) to unequivocally demonstrate the co-existence of the C-CDW and H-CDW states in thermally quenched samples. Each CDW order parameter is linked to a mirror-symmetry breaking chiral domain. The mixed-CDW state emerges from a multi-domain NC-CDW state below $T^{CD}=180$ K in cooling and $T^{WU}=210$ K in warming. The mixed-CDW low temperature state can be recovered as long as the system is not heated above $T_{I-NC}$. Scanning tunneling spectroscopy (STS) shows that the mixed-CDW is characterized by a semi-metallic in-plane density of states. At the same time, bulk resistivity remains insulating due to out-of-plane CDW stacking disorder in mixed CDW bulk samples. Our results and calculations of the free energy landscape within the Ginzburg-Landau theory demonstrate a new thermally accessible state in the phase diagram of $1T$-TaS2, enabling a new mechanism to engineer a semimetallic density of states at temperatures well above liquid nitrogen temperatures.
@@ II Results
Accessing a novel mixed-CDW state by quenching through the mirror symmetry breaking transition. In Figure 1 a, we show a characteristic [HK0] plane reciprocal space map for the C-CDW phase of $1T$-TaS2, collected after a standard cooling cycle. Red circular markers highlight the CDW reflections at $q^{C}_{1}\!=\!(\sigma^{C}_{1},\sigma^{C}_{2},l)$ and $q^{TH}_{2}\!=\!(-\sigma^{C}_{2},\sigma^{C}_{1}+\sigma^{C}_{2}+,l)$ with $\sigma^{C}_{1}\!=\!3/13$ r.l.u and $\sigma^{C}_{2}\!=\!1/13$ r.l.u Thompson et al. (1971); Wilson et al. (1975), corresponding to a single C-CDW chiral domain. Additional non-indexable reflections in the reciprocal space map are attributed to a second small crystallite in our sample. Identical reciprocal space maps were obtained across many cooling cycles and samples. We found that the C-CDW domain structure in a given sample depends on the cooling rate through $T_{\mathrm{I-NC}}$ and fast thermal quench across $T_{\mathrm{I-NC}}$ stabilizes a multi-chiral domain CDW. Fig. 1 b shows the [HK0] reciprocal space map after an in-situ quench in our bulk sample across $T_{\mathrm{I-NC}}$ down to $T\!=\!140$ K at a rate of $0.2$ K/ms near $T_{\mathrm{I-NC}}$ (See Methods for a description of the in-situ quench method and calculation of quenching rates using heat diffusion calculations). We found that thermal quenching consistently resulted in a new set of superlattice reflections of opposite chirality to the original set of satellite peaks ($\phi\!=\!-12.3^{\circ}$) (Fig. 1 b, highlighted by blue markers). The superior reciprocal space resolution of synchrotron single crystal HDRM enables us to resolve differences in the wavevector of the two sets of CDW Bragg peaks with different chirality. The in-plane wavevector magnitude $|q_{||}|\!=\!0.284(3)~{}\AA^{-1}$ of the quenched-induced satellite peaks is larger than that of the C-CDW ($|q_{||}|\!=\!0.277(3)~{}\AA^{-1}$) and is the same as the light-stabilized H-CDW state Stahl et al. (2020). This is shown in Fig. 1 c, where we display the integrated intensity across the C-CDW and H-CDW Bragg peaks (dashed lines in Fig. 1 a and b) as a function of $|q_{||}|$. Strikingly, each order parameter preferentially selects a chirality orientation, with the C-CDW domains of the mixed-CDW phase maintaining that of the as-grown samples.
@@
We also find that the c-axis periodicity differs between the C-CDW and H-CDW-like domains. Figure 1 d and e show [H0L] reciprocal space maps centered at H = 1 r.l.u. and L = 0 r.l.u. ($\delta k=0.3$ r.l.u.) before and after the thermal quench, respectively. The as-grown sample displays a broad two-peak structure, with the center of mass at integer values of L. These short-range out-of-plane correlations are characteristic of the dimerization of CDW layers in the low-temperature C-CDW phase Tanda et al. (1984). On the other hand, while retaining a dumbbell-like shape, the H-CDW-like satellite peaks shift towards $l\!=\!\pm 1/3$ indicating a collapse of the interlayer dimerization towards a triple layer CDW stacking as was observed for the light-induced H-CDW Stahl et al. (2020). The H-CDW-like satellite peaks are sharper than the C-CDW ones, reflecting a longer correlation length for the quench-induced domains $\xi\!=\!6(6)$ nm (Fig. 1 f). Similar sharpening of the out of plane CDW peaks was also observed in light-induced H-CDW $1T$-TaS2 Stahl et al. (2020). Altogether, our HDRM data in quenched samples demonstrates the coexistence of the C-CDW with a second CDW phase with the same order parameter as the H-CDW phase at much higher temperatures ($T=140$ K).
@@
The possibility of recovering the H-CDW phase as part of the stable mixed-CDW phase by keeping the samples below $T_{\mathrm{I-NC}}$ enables us to ex-situ form the CDW mixed state (see Methods) and perform additional characterization of this state. We note that the mixed CDW state is stable over months, independent of the number of cleaves, and has been reproduced in quenched samples over multiple growth batches. Detailed temperature-dependent XRD measurements (See Methods) of ex-situ quenched samples display twelve satellite reflections characteristic of the mixed-CDW state at low temperatures (Fig.2 f). For temperatures above $T^{WU}_{NC-Mixed}\!=\!210$ K during warming up, or below $T^{CD}_{NC-mixed}\!=\!195$ K during cooling down, we observe a transition into the NC-CDW. However, the NC-CDW state in quenched samples is characterized by two sets of CDW Bragg peaks with opposite chiralities. This state is shown in Fig. 2 f at $T=270$ K for ex-situ quenched samples. Equivalent HDRM data at $T=300$ K is shown in Extended Data 2. We note that the L dependence of the two NC-CDW satellite peaks is different from that of the H-CDW we observe at low temperature (see Extended data 2), ruling out a supercooled NC-CDW phase in our quenched samples Yoshida et al. (2015). The existence of two chiral domains in the NC-CDW is consistent with previous ultrafast electron diffraction experiments that demonstrated the chirality of the low-temperature C-CDW domains was inherited from the symmetry-breaking at the IC-NC transition Haupt et al. (2016) and that two NC-CDW domains were stabilized after laser excitation at room temperature Zong et al. (2018).
A detailed temperature dependence of the H-CDW-like and NC-CDW Bragg reflections are shown in Fig. 2 h and i. Upon stabilizing the mixed state at low temperatures via a quench across $T_{I-NC}$, the H-CDW peaks directly emerge from NC-CDW peaks of corresponding chirality. This temperature dependence is distinct from the metastable H-CDW phase, which is suppressed above $T=60$ K. Stahl et al. (2020); Stojchevska et al. (2014); Gerasimenko et al. (2019). We found that C-CDW and H-CDW-like Bragg peaks show the same temperature dependence (see Extended data) suggesting the coupled nature of the two order parameters in the mixed-CDW phase. The asymmetrically narrower hysteretic loop of the first-order transition, when compared to as-grown samples, $T^{CD}_{NC-C}\!=\!180$ K and $T^{WU}_{NC-C}\!=\!210$ K, in conjunction with the narrower CDW Bragg peaks along L and the temperature dependence of the [0010] structural Bragg peak shown in Fig. 2 c, suggests longer range interlayer correlations and a more 3D nature of the H-CDW phase.